Can someone help me on this problem?
I need to find the median of this set of numbers:
1, 4, 9, 15, 25, 36
Thank you,
Sher

Answer by tutorcecilia(2144)

1, 4, 9, 15, 25, 36
The median means the middle number in a set of numbers, just like the median is in the middle of the road.

First arrange the numbers in numerical order.
1, 4, 9, 15, 25, 36 (OK)


To find the middle number, start crossing-out the first and the last numbers. This leaves:
4, 9, 15, 25

Now, cross-out the next first and last numbers.
This leaves:
9, 15 are left.

Find the average of the remaining two numbers:
9 + 15 = 24/2 = 12

12 is the median or the middle of this set of numbers.